# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@ 2.2 External data
We employ several external data sets featuring measurements of BAOs, RSDs, and SN Ia to quantify the consistency of the KiDS results and infer joint constraints with the KiDS-Legacy cosmic shear data. In this section, we briefly describe the external data sets. In this work, we use publicly available likelihoods, which are implemented in the CosmoSIS standard library  [^1] https://github.com/joezuntz/cosmosis-standard-library/tree/v4.0/likelihood.
@@
We make use of BAO measurements from the first data release of the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2016, 2022) survey. In particular, DESI targets four different classes of extragalactic objects: a bright galaxy sample (BGS; Hahn et al., 2023), luminous red galaxies (LRGs; Zhou et al., 2023), emission line galaxies (ELGs; Raichoor et al., 2023), and quasars (QSOs; Chaussidon et al., 2023). Cosmological results from DESI BAO measurements were presented in Adame et al. (2025). The DESI likelihood provides BAO measurements from the four DESI sub-samples, covering a wide range of redshifts. In particular, we employ measurements from the BGS ($0.40.1<z<$ 0.4), two LRG samples ($0.60.4<z<$ 0.6 and $0.80.6<z<$ 0.8, respectively), an ELG sample ($1.61.1<z<$ 1.6), a combined LRG and ELG sample ($1.10.8<z<$ 1.1), a QSO sample ($2.10.8<z<$ 2.1), and a Lyman-$\alpha$ forest sample ($4.161.77<z<$ 4.16).
@@
As an alternative to recent BAO measurements from DESI, we employ data from the Sloan Digital Sky Survey’s (SDSS) Baryon Oscillation Spectroscopic Survey (BOSS) and Extended Baryon Oscillation Spectroscopic Survey (eBOSS). This data set provides BAO measurements as well as measurements of RSDs. The likelihoods provide constraints from the SDSS DR7 main galaxy sample (Ross et al., 2015; Howlett et al., 2015), BOSS DR12 (Alam et al., 2017), eBOSS DR16 ELGs (Tamone et al., 2020; Raichoor et al., 2021; de Mattia et al., 2021), eBOSS DR16 LRGs (Bautista et al., 2021; Gil-Marín et al., 2020), eBOSS DR16 QSOs (Neveux et al., 2020; Hou et al., 2021) and the eBOSS DR16 Lyman-$\alpha$ forest (du Mas des Bourboux et al., 2020).
@@
Additionally, we employ measurements of Type Ia supernovae from the Pantheon+ compilation (Scolnic et al., 2022). This data set consists of 1701 light curves of 1550 spectroscopically confirmed SN Ia with redshifts $z\in(0.001,2.26)$. Cosmological constraints from this data set were presented in Brout et al. (2022).
We employ cosmic shear measurements from the DES (The Dark Energy Survey Collaboration, 2005; Dark Energy Survey Collaboration et al., 2016; Flaugher et al., 2015). In particular, we make use of the ?KiDS-excised? DES data vector presented in Dark Energy Survey and Kilo-Degree Survey Collaboration et al. (2023, herafter DES+KiDS), which is based on the analysis of the DES Y3 cosmic shear measurements (Amon et al., 2022; Secco et al., 2022) and excludes 8% of DES data in the overlap region between KiDS and DES. Following the methodology of this study, we neglect the cross-covariance between the two surveys, which was shown to be sufficiently small. Furthermore, we adopt the ?$\Lambda$CDM-optimised? angular scale cuts of Amon et al. (2022) and Secco et al. (2022).
Finally, we adopt CMB measurements from Planck Collaboration et al. (2020). Here, we make use of the compressed Planck likelihood of Prince & Dunkley (2019), which approximates the $\ell<30$ temperature likelihood by two Gaussian data points and employs the Planck plik-lite TTTEEE likelihood for $\ell>30$. Following the methodology outlined in this work, we impose a Gaussian prior on the optical depth to reionisation $\tau$, which is derived from base $\Lambda$CDM parameter constraints from Planck.
@@ 3.1 Cosmic shear model
Our cosmic shear analysis pipeline is based on W25 and is implemented in the public CosmoPipe [^2] https://github.com/AngusWright/CosmoPipe/tree/KiDSLegacy_CosmicShear infrastructure. In this section, we summarise the theoretical modelling of the cosmic shear signal. We employ three summary statistics as established in the previous KiDS-1000 analysis (Asgari et al., 2021). In addition to real space shear two-point correlation functions (2PCFs), which are commonly used in cosmic shear studies, we make use of two additional summary statistics which are derived from 2PCF measurements. First, we compute complete orthogonal sets of E/B-integrals (COSEBIs; Schneider et al., 2010; Asgari et al., 2012), which provide a clean separation of E- and B-modes. This is of particular advantage since for current surveys we expect only the E modes to carry the cosmic shear signal to first order, allowing for the B modes to be used as a null test for residual systematics. Second, we employ band power spectra inferred from correlation functions (Schneider et al., 2002; Becker & Rozo, 2016; van Uitert et al., 2018). This statistic allows for an approximate separation of E- and B-modes and follows the underlying angular power spectra. In this work, we select COSEBIs as our fiducial statistic when reporting analysis results, following the choice made in W25.
In general, we model the signal for each summary statistic $S$ via a linear transformation of the cosmic shear power spectrum $C^{(ij)}_{\varepsilon\varepsilon}$:
@@
$C^{(ij)}_{\varepsilon\varepsilon}(\ell)=C^{(ij)}_{\rm GG}(\ell)+C^{(ij)}_{\rm II}(\ell)+C^{(ij)}_{\rm GI}(\ell)+C^{(ij)}_{\rm IG}(\ell)\;.$ (2)
Under the assumption of the extended Limber approximation (Kaiser, 1992; LoVerde & Afshordi, 2008; Kilbinger et al., 2017) the gravitational lensing power spectrum can be written as
@@
where $P_{\rm m,nl}$ denotes the non-linear matter power spectrum and $f_{\rm K}$, $\chi$, and $\chi_{\rm H}$ are the comoving angular diameter distance, the comoving radial distance, and the comoving horizon distance, respectively. The weak lensing kernel $W^{(i)}_{\rm G}(\chi)$ is given, for example, in eq. 2 in W25. Furthermore, the underlying galaxy sample is divided into tomographic bins using estimates of the photometric redshift, allowing for an increase in constraining power (Hu, 1999). Therefore, Eq. (3) refers to the cross cosmic shear signal between combinations of tomographic bins $i$ and $j$, where the probability distribution of comoving distances of galaxies per bin enters the window function $W^{(i)}_{\rm G}(\chi)$.
In W25 we explore a range of models of the intrinsic alignment (IA) of galaxies, finding no significant impact of the IA model on the $S_{8}$ constraints. In this work, we therefore employ the fiducial mass-dependent IA model, dubbed NLA-$M$. This model extends the non-linear linear alignment model (NLA; Bridle & King, 2007), incorporating an alignment of red, early-type galaxies, while assuming zero alignment of blue, late-type galaxies. The fraction of early-type galaxies is inferred by selecting galaxies with spectral type $T_{\rm B}<1.9$. The spectral type is inferred with bpz, which uses a set of six model templates of the spectral energy distribution (SED), ordered approximately by star formation activity, and determines the best-fitting SED by interpolating between templates. The cut on $T_{\rm B}$ selects galaxies with contributions of an elliptical galaxy spectrum. We model the IA power spectrum of red galaxies as a power law dependent on the average halo mass within a tomographic bin (see eq. 11 in W25, ). In this model, we employ two nuisance parameters, $A_{\rm IA}$ and $\beta$, which parameterise the IA amplitude and the slope of the IA mass scaling, respectively. We adopt the joint posterior on $A_{\rm IA}$ and $\beta$ from Fortuna et al. (2025) as prior in our analysis, which we approximate by a bivariate Gaussian distribution. Additionally, we employ a multivariate Gaussian prior on the halo mass per tomographic bin. For a detailed description of the IA model, we refer the reader to section 2.3.4 and appendix B in W25.
@@ 3.2 Consistency metrics
Table 1: Model parameters and their priors.
@@
Cosmological $\omega_{\rm cdm}$ $\mathcal{U}(0.051,0.255)$ Yes Reduced cold dark matter density
@@
Nuisance $\log_{10}T_{\rm AGN}$ $\mathcal{U}(7.3,8.3)$ Yes Baryon feedback parameter
$A_{\rm IA}$ $\mathcal{N}(5.74,\mathbf{C}_{A_{\rm IA},\beta})$ No / red-blue split only Amplitude of intrinsic galaxy alignments for red galaxies
@@
$\delta_{{\rm z},i}$ $\mathcal{N}(\bm{\mu},\mathbf{C}_{z})$ No / red-blue split only Shift of the mean of the redshift distribution per tomographic bin $i$
 [^3] The first two columns provide parameter names and types of the sampling parameters. The third column lists the adopted prior with uniform priors denoted by $\mathcal{U}$ and Gaussian priors denoted by $\mathcal{N}$. The fourth column indicates whether or not the parameter is duplicated in the split cosmological analysis. The fifth column provides a brief parameter description. The priors on $A_{\rm IA}$ and $\beta$ are centred at the mean of the posterior reported by Fortuna et al. (2025) including a correlated prior with correlation coefficient $r=-0.59$. The prior on $M_{\rm i}$ is inferred from halo mass measurements as discussed in appendix B of W25. The $\delta_{\rm z}$ priors are centred at the mean of the shift of the redshift distribution and correlated through a covariance matrix estimated following the methodology of Wright et al. (2025a).
To assess the internal consistency of the KiDS-Legacy data set, we follow the methodology of Köhlinger et al. (2019) and subdivide the data in many ways before analysing the subsets jointly. In particular, we apply splits at the data vector level by redshift and angular scale and at the catalogue level by spatial region and colour. Additionally, we construct a joint data vector of different summary statistics. For each split, we analyse the cosmic shear data with two modelling setups:
@@
    Fiducial cosmological model: one set of parameters models the full data set.
@@
For splits at the data vector-level the first setup is equivalent to the fiducial cosmic shear analysis setup of W25. For splits at the catalogue level and splits by summary statistic we construct a different data vector, whose information content may differ from the fiducial data vector: for example, the red-blue split excludes shape correlations between red and blue galaxies. The second modelling setup features two independent sets of cosmological parameters, which allow us to assess whether or not different subsets of the data prefer different cosmologies. A list of cosmological and nuisance parameters is given in Table 1. The cosmic shear analysis features a number of nuisance parameters which are marginalised over, assuming Gaussian or top-hat priors. Since the posterior distribution of the nuisance parameters is entirely driven by the prior (see W25, ) we do not duplicate these parameters in the split cosmological analysis, but instead keep them shared between both data subsets. An exception is the colour-based split, as discussed in Sect. 4.2.2.
In practice, we conduct a likelihood analysis for each data split employing both the fiducial model and the split cosmological model. We then evaluate various consistency metrics, testing if the split cosmological model is preferred over the fiducial model. There exists a variety of statistical tools allowing for model comparison and an estimation of the significance of the preference of a specific model. These can be grouped into techniques that compress the full likelihood or posterior into a single summary statistic, parameter-space methods that focus on differences in single or multiple model parameters, and methods that quantify differences in data vector space. In this work, we perform three tiers of consistency tests, following the nomenclature of Köhlinger et al. (2019). We note that only the tier 1 test requires a cosmological analysis with the fiducial model since it performs a model comparison. The remaining tiers focus on the analysis with the split cosmological model in order to quantify the consistency between data subsets.
@@ 3.2.1 Tier 1: Evidence-based metric
The first tier of consistency tests includes tests compressing the full likelihood or posterior into a single summary statistic. A widely used example of such a metric is the Bayes ratio, which in logarithmic form is given by the difference between Bayesian evidences
@@
where $\mathcal{Z}_{\rm fiducial/split}$ denotes the Bayesian evidence of the two models which, in general, can be computed by integrating the product of the likelihood $\mathcal{L}$ and the prior $\pi$
@@
over the parameters ${\bm{\theta}}$. We note that for two independent data sets A and B, the evidence for the split model simplifies to $\mathcal{Z}_{\rm split}=\mathcal{Z}_{\rm A}\mathcal{Z}_{\rm B}$. However, this equality does not hold for correlated data sets, i.e. for most data splits considered in this work. Therefore, the evidence needs to be computed from the joint posterior distribution of the two subsets, modelled with two sets of parameters, and taking the cross-correlation between subsets into account.
In general, values of $\log_{10}R>0$ corresponds to preference for the fiducial model while $\log_{10}R<0$ indicates preference for the split model. The Bayes ratio is usually interpreted using Jeffreys’ scale (Jeffreys, 1939), which provides limits for the degree of preference for a specific model. This scale associates values of $|\log_{10}R|>[\frac{1}{2},1,2]$ with ?substantial?, ?strong?, and ?decisive? preference for the specific model, respectively, but this lacks a clear motivation and there is no consensus on when to report tension between models. Additionally, the Bayes ratio suffers from a prior dependence which is suboptimal when using wide, uninformative priors. This makes the Bayes ratio particularly suboptimal for the analysis presented in this work, since it involves a duplication of the parameter space and the corresponding prior volume.
To circumvent these issues, Handley & Lemos (2019b) proposed the so-called suspiciousness $S$:
$\ln S=\ln R-\ln I\;,$ (6)
where the information ratio $\ln I=\mathcal{D}_{\rm split}-\mathcal{D}_{\rm fiducial}$ is defined through the Kullback-Leibler divergence $\mathcal{D}$ (Kullback & Leibler, 1951), which quantifies the information gain between the prior and the posterior. The suspiciousness is designed to remove the effect of the prior volume from the Bayes ratio. As discussed in Handley & Lemos (2019b), it is insensitive to the choice of prior as long as the prior does not affect the shape of the posterior distribution. Under the assumption of Gaussian posteriors, a tension probability can be identified via the quantity $d-2\ln S$. Here, $d$ denotes the difference between the effective number of constrained parameters by the two models
@@
In a prior-informed analysis with correlated sampling parameters, such as the cosmic shear analysis in this work, this quantity is smaller than the number of free parameters and is non-trivial to determine. There exist several estimators, such as the Bayesian model complexity (BMC, Spiegelhalter et al., 2002) and the Bayesian model dimensionality (BMD, Handley & Lemos, 2019a). However, Joachimi et al. (2021) report that commonly used dimensionality measures in general are biased estimators of the effective number of parameters. As an alternative, they proposed an estimation via $\chi^{2}$ minimisation of a set of mock data vectors, which on average was found to reproduce an unbiased estimate of the true value of $N_{\Theta}$. Therefore, we adopt this strategy as the fiducial method in the analysis. For comparison, we additionally compute the BMD given by
@@
where $\langle\rangle_{P}$ denotes the average over the posterior. This quantity can be directly obtained as a byproduct of common posterior sampling algorithms, such as MCMC or nested sampling, and therefore is available at no additional computational cost. For Gaussian posteriors, the tension probability inferred from the suspiciousness statistic is then determined by

### v2

@@ 2.2 External data
We employed several external datasets featuring measurements of BAOs, RSDs, and SN Ia to quantify the consistency of the KiDS results and infer joint constraints with the KiDS-Legacy cosmic shear data. In this section, we briefly describe the external datasets. In this work, we use publicly available likelihoods, which are implemented in the CosmoSIS standard library  [^1] https://github.com/joezuntz/cosmosis-standard-library/tree/v4.0/likelihood.
@@
We made use of BAO measurements from the first data release of the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2016, 2022) survey. In particular, DESI targets four different classes of extragalactic objects: a bright galaxy sample (Hahn et al., 2023), luminous red galaxies (LRGs; Zhou et al., 2023), emission line galaxies (ELGs; Raichoor et al., 2023), and quasi-stellar objects (QSOs; Chaussidon et al., 2023). Cosmological results from DESI BAO measurements were presented in Adame et al. (2025). The DESI likelihood provides BAO measurements from the four DESI sub-samples, covering a wide range of redshifts. In particular, we employ measurements from the bright galaxy sample ($0.1<z<0.4$), two LRG samples ($0.4<z<0.6$ and $0.6<z<0.8$, respectively), an ELG sample ($1.1<z<1.6$), a combined LRG and ELG sample ($0.8<z<1.1$), a QSO sample ($0.8<z<2.1$), and a Lyman-$\alpha$ forest sample ($1.77<z<4.16$).
@@
As an alternative to recent BAO measurements from DESI, we employ data from the Sloan Digital Sky Survey’s (SDSS) Baryon Oscillation Spectroscopic Survey (BOSS) and Extended Baryon Oscillation Spectroscopic Survey (eBOSS). This dataset provides BAO measurements as well as measurements of RSDs. The likelihoods provide constraints from the SDSS DR7 main galaxy sample (Ross et al., 2015; Howlett et al., 2015), BOSS DR12 (Alam et al., 2017), eBOSS DR16 ELGs (Tamone et al., 2020; Raichoor et al., 2021; de Mattia et al., 2021), eBOSS DR16 LRGs (Bautista et al., 2021; Gil-Marín et al., 2020), eBOSS DR16 QSOs (Neveux et al., 2020; Hou et al., 2021), and the eBOSS DR16 Lyman-$\alpha$ forest (du Mas des Bourboux et al., 2020).
@@
Additionally, we employed measurements of SN Ia from the Pantheon+ compilation (Scolnic et al., 2022). This dataset consists of 1701 light curves of 1550 spectroscopically confirmed SN Ia with redshifts $z\in(0.001,2.26)$. Cosmological constraints from this dataset were presented in Brout et al. (2022).
[…]
@@ 3.1 Cosmic shear model
Our cosmic shear analysis pipeline is based on W25 and is implemented in the public CosmoPipe [^2] https://github.com/AngusWright/CosmoPipe/tree/KiDSLegacy_CosmicShear infrastructure. In this section, we summarise the theoretical modelling of the cosmic shear signal. We employed three summary statistics as established in the previous KiDS-1000 analysis (Asgari et al., 2021). In addition to real space shear two-point correlation functions (2PCFs), which are commonly used in cosmic shear studies, we made use of two additional summary statistics which are derived from 2PCF measurements. First, we computed complete orthogonal sets of E/B-integrals (COSEBIs; Schneider et al., 2010; Asgari et al., 2012), which provide a clean separation of E and B modes. This is of particular advantage since, for current surveys, we expect only the E modes to carry the cosmic shear signal to first order, allowing for the B modes to be used as a null test for residual systematics. Second, we employed band power spectra inferred from correlation functions (Schneider et al., 2002; Becker & Rozo, 2016; van Uitert et al., 2018). This statistic enables an approximate separation of E and B modes and follows the underlying angular power spectra. In this work, we used COSEBIs as our fiducial statistic when reporting analysis results, following the choice made in W25.
In general, we model the signal for each summary statistic $S$ via a linear transformation of the cosmic shear power spectrum, $C^{(ij)}_{\varepsilon\varepsilon}$:
@@
[…]
Under the assumption of the extended Limber approximation (Kaiser, 1992; LoVerde & Afshordi, 2008; Kilbinger et al., 2017), the gravitational lensing power spectrum can be written as
@@
where $P_{\rm m,nl}$ denotes the non-linear matter power spectrum and $f_{\rm K}$, $\chi$, and $\chi_{\rm H}$ are the comoving angular diameter distance, the comoving radial distance, and the comoving horizon distance, respectively. The weak lensing kernel $W^{(i)}_{\rm G}(\chi)$ is given, for example, in eq. 2 in W25. Furthermore, the underlying galaxy sample is divided into tomographic bins using estimates of the photometric redshift, allowing for an increase in constraining power (Hu, 1999). Therefore, Eq. (3) refers to the cross cosmic shear signal between combinations of tomographic bins $i$ and $j$, where the probability distribution of comoving distances of galaxies per bin enters the window function, $W^{(i)}_{\rm G}(\chi)$.
In W25, we explored a range of models of the intrinsic alignment (IA) of galaxies, finding no significant impact of the IA model on the $S_{8}$ constraints. In the present work, we therefore employed the fiducial mass-dependent IA model, dubbed NLA-$M$. This model extends the non-linear linear alignment (NLA) model (Bridle & King, 2007), incorporating an alignment of red, early-type galaxies, while assuming zero alignment of blue, late-type galaxies. The fraction of early-type galaxies is inferred by selecting galaxies with spectral type $T_{\rm B}<1.9$. The spectral type is inferred with bpz, which uses a set of six model templates of the spectral energy distribution ordered approximately based on the star formation activity and determines the best-fitting spectral energy distribution by interpolating between templates. The cut on $T_{\rm B}$ selects galaxies with contributions of an elliptical galaxy spectrum. We modelled the IA power spectrum of red galaxies as a power law dependent on the average halo mass within a tomographic bin (see eq. 11 in W25, ). In this model, we employed two nuisance parameters, $A_{\rm IA}$ and $\beta$, which parameterise the IA amplitude and the slope of the IA mass scaling, respectively. We adopted the joint posterior on $A_{\rm IA}$ and $\beta$ from Fortuna et al. (2025) as a prior in our analysis, which we approximated by a bivariate Gaussian distribution. Additionally, we employed a multivariate Gaussian prior on the halo mass per tomographic bin. For a detailed description of the IA model, we refer to Section 2.3.4 and appendix B in W25.
@@ 3.2 Consistency metrics
Table 1: Model parameters and their priors. [^3] The first two columns provide parameter names and types of the sampling parameters. The third column lists the adopted prior with uniform priors denoted by $\mathcal{U}$ and Gaussian priors denoted by $\mathcal{N}$. The fourth column indicates whether or not the parameter is duplicated in the split cosmological analysis. The fifth column provides a brief parameter description. The priors on $A_{\rm IA}$ and $\beta$ are centred at the mean of the posterior reported by Fortuna et al. (2025) including a correlated prior with correlation coefficient r=0.59r=-0.59. The prior on $M_{\rm i}$ is inferred from halo mass measurements as discussed in appendix B in W25. The $\delta_{\rm z}$ priors are centred at the mean of the shift of the redshift distribution and correlated through a covariance matrix estimated following the methodology of Wright et al. (2025a).
@@
Cosmological
$\omega_{\rm cdm}$ $\mathcal{U}(0.051,0.255)$ Yes Reduced cold dark matter density
@@
Nuisance
$\log_{10}T_{\rm AGN}$ $\mathcal{U}(7.3,8.3)$ Yes Baryon feedback parameter
$A_{\rm IA}$ $\mathcal{N}(5.74,\mathbf{C}_{A_{\rm IA},\beta})$ No + red-blue split only Amplitude of intrinsic galaxy alignments for red galaxies
@@
$\delta_{{\rm z},i}$ $\mathcal{N}(\bm{\mu},\mathbf{C}_{z})$ No + red-blue split only Shift of the mean of the redshift distribution per tomographic bin $i$
To assess the internal consistency of the KiDS-Legacy dataset, we followed the methodology of Köhlinger et al. (2019) and subdivided the data in many ways before analysing the subsets jointly. In particular, we applied splits at the data vector level by redshift and angular scale and at the catalogue level by spatial region and colour. Additionally, we constructed a joint data vector of different summary statistics. For each split, we analysed the cosmic shear data with two modelling setups:
@@
    Fiducial cosmological model: one set of parameters models the full dataset;
@@
For splits at the data vector-level, the first setup is equivalent to the fiducial cosmic shear analysis setup of W25. For splits at the catalogue level and splits by summary statistic, we constructed a different data vector, whose information content may differ from the fiducial data vector; for example, the red-blue split excludes shape correlations between red and blue galaxies. The second modelling setup features two independent sets of cosmological parameters, which allow us to assess whether or not different subsets of the data prefer different cosmologies. A list of cosmological and nuisance parameters is given in Table 1. The cosmic shear analysis features a number of nuisance parameters which are marginalised over, assuming Gaussian or top-hat priors. Since the posterior distribution of the nuisance parameters is entirely driven by the prior (see W25, ) we did not duplicate these parameters in the split cosmological analysis; instead, we kept them shared between both data subsets. An exception is the colour-based split, as discussed in Sect. 4.2.2.
In practice, we conducted a likelihood analysis for each data split employing both the fiducial model and the split cosmological model. We then evaluated various consistency metrics, testing whether the split cosmological model is preferred over the fiducial model. There are a variety of statistical tools available that allow for a model comparison and an estimation of the significance of the preference of a specific model. These can be grouped into techniques that compress the full likelihood or posterior into a single summary statistic, parameter-space methods that focus on differences in single or multiple model parameters, as well as methods that quantify differences in data vector space. In this work, we performed three tiers of consistency tests, following the nomenclature of Köhlinger et al. (2019). We note that only the tier 1 test requires a cosmological analysis with the fiducial model since it performs a model comparison. The remaining tiers focus on the analysis with the split cosmological model to quantify the consistency between data subsets.
@@ 3.2.1 Tier 1: Evidence-based metric
The first tier of consistency tests includes tests compressing the full likelihood or posterior into a single summary statistic. A widely used example of such a metric is the Bayes ratio, which in logarithmic form is given by the difference between Bayesian evidence,
@@
where $\mathcal{Z}_{\rm fiducial/split}$ denotes the Bayesian evidence of the two models, which can generally be computed by integrating the product of the likelihood, $\mathcal{L}$, and the prior, $\pi$,
@@
over the parameters represented by $\bm{\theta}$}. We note that for two independent datasets, A and B, the evidence for the split model simplifies to $\mathcal{Z}_{\rm split}=\mathcal{Z}_{\rm A}\mathcal{Z}_{\rm B}$. However, this equality does not hold for correlated datasets, namely, for most data splits considered in this work. Therefore, the evidence needs to be computed from the joint posterior distribution of the two subsets, modelled with two sets of parameters, and taking the cross-correlation between subsets into account.
In general, values of $\log_{10}R>0$ correspond to preference for the fiducial model while $\log_{10}R<0$ indicates preference for the split model. The Bayes ratio is usually interpreted using Jeffreys’ scale (Jeffreys, 1939), which provides limits for the degree of preference for a specific model. This scale associates values of $|\log_{10}R|>[\frac{1}{2},1,2]$ with ?substantial?, ?strong?, and ?decisive? preference for the specific model, respectively, but this lacks a clear motivation and there is no consensus on when to report tension between models. Additionally, the Bayes ratio suffers from a prior dependence which is suboptimal when using wide, uninformative priors. This makes the Bayes ratio particularly suboptimal for the analysis presented in this work, since it involves a duplication of the parameter space and the corresponding prior volume.
To circumvent these issues, Handley & Lemos (2019b) proposed the so-called suspiciousness parameter, $S$, expressed as
$\ln S=\ln R-\ln I\;.$ (6)
Here, the information ratio $\ln I=\mathcal{D}_{\rm split}-\mathcal{D}_{\rm fiducial}$ is defined through the Kullback-Leibler divergence $\mathcal{D}$ (Kullback & Leibler, 1951), which quantifies the information gain between the prior and the posterior. The suspiciousness is designed to remove the effect of the prior volume from the Bayes ratio. As discussed in Handley & Lemos (2019b), it is insensitive to the choice of prior as long as the prior does not affect the shape of the posterior distribution. Under the assumption of Gaussian posteriors, a tension probability can be identified via the quantity $d-2\ln S$. Here, $d$ denotes the difference between the effective number of constrained parameters by the two models
@@
In a prior-informed analysis with correlated sampling parameters, such as the cosmic shear analysis in this work, this quantity is smaller than the number of free parameters and is non-trivial to determine. There exist several estimators, such as the Bayesian model complexity Spiegelhalter et al. (2002) and the Bayesian model dimensionality (BMD; Handley & Lemos, 2019a). However, Joachimi et al. (2021) reported that commonly used dimensionality measures in general are biased estimators of the effective number of parameters. As an alternative, they proposed an estimation via $\chi^{2}$ minimisation of a set of mock data vectors, which was generally found to reproduce an unbiased estimate of the true value of $N_{\Theta}$. Therefore, we adopted this strategy as the fiducial method in the analysis. For comparison, we additionally computed the BMD via
@@
where $\langle\rangle_{P}$ denotes the average over the posterior. This quantity can be directly obtained as a byproduct of common posterior sampling algorithms, such as Markov chain Monte Carlo or nested sampling; therefore, it is available at no additional computational cost. For Gaussian posteriors, the tension probability inferred from the suspiciousness statistic is then determined by